# Testing the Fifth Force on Lepton Spins through Neutrino Oscillations

### v1

@@ B.1 Reactor Long-baselineexperiments
where $i$ is the index of the energy bin, $k$ represent the fission isotopics $({}^{235}\!\rm{U},\allowbreak{}^{238}\!\rm{U},\allowbreak{}^{239}\!\rm{Pu},\allowbreak{}^{241}\!\rm{Pu})$, $f_{k}$ is the contribution of each fission isotopic, taken to be $(0.567,\allowbreak 0.078,\allowbreak 0.298,\allowbreak 0.057)$ [62]. $\Phi_{k}$ is the anti-neutrino flux from the fission of each isotopic, as given in Ref. [64]. Our results show that the calculated oscillation probabilities agree well with those in Ref. [62]. Based on this consistency, we use our calculated survival probabilities to place constraints on the fifth force model using Eq. (9). The observed survival probabilities and uncertainties are taken directly from Ref. [62].
B.2 Reactor Medium-baseline experiments
@@
Going through three stages with different amount of ADs, Ref. [68] provides the observed survival probability $P(\bar{\nu}_{e}\rightarrow\bar{\nu}_{e})$ corresponding to $L_{\rm eff}/\langle E_{\bar{\nu}_{e}}\rangle$, where $\langle E_{\bar{\nu}_{e}}\rangle$ is the mean antineutrino energy, and the effective baseline $L_{\rm eff}$ represents the effective baseline. The effective baselines for EH1, EH2, and EH3 are 500 m, 500 m, and 1650 m, respectively [68]. Using the survival probabilities $P(\bar{\nu}_{e}\rightarrow\bar{\nu}_{e})$ observed at different energy points by the detectors in EH1, EH2, and EH3, we can place constraints on the fifth force model, as described using Eq. (10).

### v2

@@ IX.1 Reactor Long-baseline experiments
where $i$ is the index of the energy bin, $k$ represent the fission isotopics $({}^{235}\!\rm{U},\allowbreak{}^{238}\!\rm{U},\allowbreak{}^{239}\!\rm{Pu},\allowbreak{}^{241}\!\rm{Pu})$, $f_{k}$ is the contribution of each fission isotopic, taken to be $(0.567,\allowbreak 0.078,\allowbreak 0.298,\allowbreak 0.057)$ [63]. $\Phi_{k}$ is the anti-neutrino flux from the fission of each isotopic, as given in Ref. [65]. Our results show that the calculated oscillation probabilities agree well with those in Ref. [63]. Based on this consistency, we use our calculated survival probabilities to place constraints on the fifth force model using Eq. (10) in the main text. The observed survival probabilities and uncertainties are taken directly from Ref.[63].
IX.2 Reactor Medium-baseline experiments
@@
Going through three stages with different amount of ADs, Ref. [69] provides the observed survival probability $P(\bar{\nu}_{e}\rightarrow\bar{\nu}_{e})$ corresponding to $L_{\rm eff}/\langle E_{\bar{\nu}_{e}}\rangle$, where $\langle E_{\bar{\nu}_{e}}\rangle$ is the mean antineutrino energy, and the effective baseline $L_{\rm eff}$ represents the effective baseline. The effective baselines for EH1, EH2, and EH3 are 500 m, 500 m, and 1650 m, respectively [69]. Using the survival probabilities $P(\bar{\nu}_{e}\rightarrow\bar{\nu}_{e})$ observed at different energy points by the detectors in EH1, EH2, and EH3, we can place constraints on the fifth force model, as described using Eq. (10) in the main text.